OpenStudy (anonymous):

Solve the equation: –2a – 11a = –5 and check

OpenStudy (anonymous):

-2a+11a= 9a=-5 then bring over the 9 to the other side a=-5/9

OpenStudy (anonymous):

-2a-11a=-5 -13a=-5 a=5/13

OpenStudy (anonymous):

@kungfupichu -2a-11a=-2a + (-11a) which gets you -13a as @hasteeerfani00 said.... so then you get -13a=-5 and so forth... and to check, just plug all values of a with 5/13 and if both sides are the same... then it is correct!

OpenStudy (anonymous):

thxx

OpenStudy (anonymous):

np :)